# LLMs for Generalizable Language-Conditioned Policy Learning under Minimal Data Requirements

### v1

@@ 3.1.1 State abstraction
where $\hat{P}^{g}$ is the empirical state transition function based on the abstract datasets ${\mathcal{D}}^{g}$, obtained during Q-learning in step 2. We then fine-tune a pre-trained large language model on ${\mathcal{D}}^{SFT}$ using the standard next-word prediction objective. We integrate description of the goal $g$ and the state $s^{g}_{0}$ into a prompt and set the sequence $[a_{0}^{*,g},\ldots,a_{n_{g}}^{*,g}]$ as the expected completion. We expect that the fine-tuned language model combined with the state abstraction function $\mathrm{LLM}^{abstrct}$ can effectively act as a proxy for the general, goal-conditioned policy $\pi^{*}$ from equation (2), generalizing to any new goal $g\notin{\mathcal{G}}^{tr}$ and previously unobserved low-level state $x\in{\mathcal{X}}$.
@@ 4 Related Work
LLMs for decision making. There is growing interest in using general-purpose LLMs directly as decision-making agents (Yao et al., 2023). Various prompting techniques, such as chain of thought (Wei et al., 2023) and self-reflection (Ji et al., 2023), have been developed to enhance LLMs’ abilities in long-term planning tasks. However, as demonstrated in previous works (Szot et al., 2024; Finn, 2024), prompting techniques alone are insufficient for solving complex decision-making tasks in dynamic environments. To effectively utilize the knowledge embedded in LLMs for RL problems, these models must be grounded in the dynamics of the environment. This grounding can be achieved either through in-context learning (Wang et al., 2023; Wu et al., 2023) or fine-tuning (Carta et al., 2023; Tan et al., 2024; Brohan et al., 2023a). A key limitation of in-context learning is its restricted window size. In this work, we focus on fine-tuning; however, unlike prior studies, we significantly reduce the requirements on input data for fine-tuning the decision-making agent.
LLMs as data enhancers. To apply conventional RL methods in search of optimal goal-conditioned policies, we must augment our dataset of state-action transitions with goal-dependent rewards. This process, known as hindsight labeling, has traditionally been performed manually by human annotators or through learning the reward function from expert demonstrations (Ziebart et al., 2008; Fu et al., 2018; Bahdanau et al., 2018). Recent studies, however, have demonstrated that task-specific rewards can be effectively generated using pre-trained LLMs (Yu et al., 2023b; Ma et al., 2023; Xie et al., 2023). While successful, most LLM-based approaches rely on iterative prompting strategies, which are costly in terms of LLM calls. Our approach to hindsight labeling reduces this cost by approximating the LLM-induced reward function with a lightweight neural network. Furthermore, we assign rewards in abstracted state spaces, significantly reducing the number of states to be labeled. Similar to the work of Peng et al. (2023), our state abstraction function uses the language command to guide the elimination of irrelevant state features.
Language-conditioned RL. Numerous previous studies have explored learning language-conditioned policies by assuming access to ground-truth environment rewards (Jiang et al., 2019; Co-Reyes et al., 2018), real-time experimentation (Fu et al., 2018; Bahdanau et al., 2018; Mirchandani et al., 2021), or expert demonstrations paired with language annotations (Stepputtis et al., 2020; Lynch & Sermanet, 2021; Xiao et al., 2023; Brohan et al., 2023b; a). In contrast, our approach aims to learn language-conditioned policies from entirely offline datasets, which may be highly suboptimal and which are unlabeled, with no environment- or human-provided reward signals. Regarding policy evaluation, much of the prior work in language-conditioned RL and IL tests agents on new language commands synonymous with those seen during training (Lynch & Sermanet, 2021; Nair et al., 2022). Similar to the works of (Xiao et al., 2023; Brohan et al., 2023a; Shridhar et al., 2021a; Jang et al., 2022), our focus lies on novel instructions corresponding to previously unsolved goals.
Refer to Appendix A for an extended discussion of the related work.
@@ 5 Experiments
Experimental Setup. For the experiments we require a controlled environment where a wide variety of distinct goals can be specified, and where the semantics of the states and actions are easily interpretable by an LLM. Thus, we choose the BabyAI environment (Chevalier-Boisvert et al., 2018), a grid world platform for instruction following where an agent receives natural language goal instructions such as: “Go to the tile (3,2)”, “Pick up the blue key” or “Look behind the green locked door”. A detailed discussion on the environment choice can be found in Appendix A. The grids can consist of multiple rooms connected by open or locked doors and different distractor objects that the agent can interact with (boxes, keys, balls, etc.). The action space $\mathcal{A}$ consists of several navigation primitives (forward, pickup, etc.). In our setup, we assume full observability of the original state space. Each state can be represented as a long list of features and their coordinates (see Appendix B.2 for example state representations in a text format).
5.1 Q1: Online Evaluation: Generalization Benchmark
Setup. We choose the collection of Synth environments from BabyAI as the main test bed for TEDUO. All environments are constructed as a 22x22 grid and containing 9 rooms. They differ in the type, position, and color of the distractors. The tasks include goals such as “go to the {color} {object}”, “pick up the {color} {object}”, or “put the {color} {object} next to the {color} {object}”. We use a list of 500 goals as ${\mathcal{G}}^{tr}$. The set of testing goals contains 100 goals that are semantically distinct from those in ${\mathcal{G}}^{tr}$. We also augment the set of testing goals by asking GPT-4 to paraphrase the original commands provided by BabyAI. We train a Llama-3-8B model with TEDUO based on a dataset ${\mathcal{D}}$ containing 800k non-unique state-action-state triplets generated according to a policy that is a random mixture of default policies from BabyAI (see Appendix B.1 for details).
Baselines. We compare our fine-tuned Llama-3-8B agent with non-fine-tuned LLMs: Llama-3-8B and Llama-3-70B using a) vanilla and b) chain-of-thought prompting Wei et al. (2023) with additional demonstrations provided in-context (in-context+CoT). The latter integrates expert demonstrations generated during step 2 of TEDUO to test the in-context learning ability of the LLM. Following recent works Mezghani et al. (2023); Li et al. (2022); Cao et al. (2023), we also compare against BabyAI-IL-bot, the baseline proposed by the authors of BabyAI (Chevalier-Boisvert et al., 2018), which is the combination of a GRU to encode the instruction, CNN+FILM layers to encode the grid and an LSTM memory. We train this method via imitation learning on the policy generated by TEDUO, steps 1&2. Implementation details of the baselines can be found in Appendix B.6.
Table 1: Online evaluation of generalization performance. Results averaged over 400 $(g,s_{0}^{g})$ pairs. Standard error in brackets.
@@ 5.1 Q1: Online Evaluation: Generalization Benchmark
  • Prior knowledge of LLMs is insufficient. We find that non-fine-tuned LLMs, irrespective of their parameter count or prompting method struggle in solving tasks from the BabyAI environments. Low success rate and high invalid action ratios indicate the inability of LLMs to understand the dynamics of the environment. Common failures include only using the action “move forward” without considering the agent’s direction or attempting final actions (e.g. door opening) without first navigating to the correct location, The observed poor performance, underscores the need for developing data-efficient methods of distilling knowledge of the environment-dynamics into LLMs. Our fine-tunning strategy brings the success rate of the Llama-3-8B language model from 17% to 65% in its training setting and 45% for the in-the-wild setting.
  • Generalization to new environments and goals. We further look at the generalization abilities of our fine-tuned TEDUO-Llama-3-8B model to new environments and goals. When the LLM is tested on new environments unseen during training, a performance drop of 12% is observed, significantly lower than BabyAI-IL-bot baseline with a drop of 24%. This difference can be explained by the RL baseline’s overfitting due to the limited offline training data, while TEDUO-Llama-3-8B benefits from the zero-shot capabilities of the pretrained LLM. This effect is even more pronounced with new goals, where TEDUO experiences only an 8% decrease in success rate, compared to a 40% drop for the BabyAI-IL-bot. Overall, TEDUO achieves nearly three times better performance than the RL baseline when generalizing to both new natural language commands and environments. We analyze success rates per goal type in Appendix D.1.
@@ 5.2 Q2: Online Evaluation: Ablation Study
Setup. Using the same experimental setup as in the previous section, we compare our full fine-tuning pipeline with its ablations. After obtaining the abstract datasets ${\mathcal{D}}^{g}$ with the first step of TEDUO, we generate goal-conditioned policies with naive behavioral cloning (step 1 + GCBC). We also compare our fine-tuned Llama-3-8B against the performance of the GCRL policies obtained with offline Q-learning in step 2. Note, neither of GCBC nor GCRL can generalize to new, previously unseen language commands. Therefore, in this study, we are only looking at performance on goals from ${\mathcal{G}}^{tr}$. Ablation of the abstraction function is delayed to the next section.
Table 2: Ablation study. Results averaged over 400 $(g,s_{0}^{g})$ pairs. Standard error in brackets.
@@
Results. The results of GCBC and GCRL can be seen as ablations of our pipeline. We first note that the success rate of naive behavioural cloning is low, indicating low fidelity of the data collection policy and highlighting the need for incorporating offline policy-learning methods. Moreover, the significantly improved performance of the Q-learning policies (GCRL) validates the effectiveness of the first two steps within our pipeline. Thus, the synthetically constructed abstract MDPs are meaningful offline constructs that yield, given the data available, optimal policies effective during online testing. Finally, the improved performance of our fine-tuned Llama-3-8B over the Q-learning/GCRL policies on training goals and environments confirms the importance of the third step of our method and suggests that the ungrounded, prior knowledge of large language models improves generalization to new previously unseen states.
@@ 5.3 Q3: Learning and exploiting core skills
The aim of the first part of our experiments is to investigate the generalization abilities of the LLM fine-tuned with our pipeline. We wish to investigate if by learning the optimal policies for diverse goals and environments, the LLM can integrate the core skills required to achieve these goals and how such skills can be transferred across tasks. We also investigate the aspect of skill compositionality. Does the prior knowledge of the LLM, now grounded in the environment dynamics, suffice to compose together learned skills to solve novel tasks?
@@
Setup. We are working with the following three types of illustrative environments:
[Uncaptioned image] Type A: A grid with 2 rooms, an open door and a box. The language commands are of two types: “go to the tile (x,y)” and “pick up the {color} box”.
[Uncaptioned image] Type B: A grid with 2 rooms and a closed door. The language commands are the types: “go to the tile (x,y)” and “open the {color} door”. The agent and the object are in different rooms, so the agent must pick up the key and open the door first.
[Uncaptioned image] Type C: A grid with 2 rooms, a closed door and a box. The language commands are the types: “pick up the {color} box”. The agent and the object are in different rooms, so the agent must open the door first and then pick up the box.
The position and color of the door and box vary across different instantiations of the environments. We use type A and B environments for training and type C environments for testing. We note that tasks from type C environments require the internalization of three core skills: moving to a given location, opening a door, picking up a box. The skill of moving to a location can be obtained from both environments A and B, but the skill of picking up a box or opening the door can only be obtained from one of the environments, A or B, respectively. This setup allows us to investigate the transferability of learned skills across environments and their compositionality.
Table 3: Performance on test tasks from type C environments. TEDUO A and TEDUO B have been trained in only one environment whereas TEDUO A&B has been trained in both.
@@
Results. Table 3 demonstrates that the LLM fine-tuned on tasks from both Type A and B environments achieves a 60% success rate, compared to the non-fine-tuned baseline, which fails entirely. Although the environment is simpler, baseline performance is lower than in Table 1 due to shorter maximum episode length (20 vs. 500). The agent trained only in Environment A (TEDUO A) reaches a 99% success rate in tasks without closed doors (Type A grids), but consistently fails when the goal is behind a door. Similarly, TEDUO B achieves an 81% success rate in new grids from Type B but cannot generalize to Type C. The observation that TEDUO A&B can generalize to a new environment C that requires a combination of both skills independently seen during training indicates that the fine-tuned LLM does not merely memorize optimal trajectories for individual tasks. Instead, it learns core, generalizable abilities that can be combined to address novel tasks in new settings. This result contrasts with the failure of an LLM trained in only one environment, emphasizing the significance of multi-skill learning for successful generalization, which our framework enables.
@@
One of the core skills required to successfully solve tasks from the BabyAI environments is to identify whether the agent at its given location is facing an object or a wall, or it is free to move forwards. This section provides additional insights into the behaviour and internal representation of states of the LLM fine-tuned with TEDUO in comparison to a base LLM.
@@
Setup. We operate within the Synth BabyAI environments as in the main evaluation benchmark and generate a dataset consisting of 10 random goals and 512 states per each goal. We embed each goal-state pair into our prompt template for eliciting actions and fine-tuning the language models and pass them through both the base and fine-tuned Llama-3-8B from experiments 5.1 and 5.2. We record the logprobabilities of the tokens [0, 1, …, 6] as well as the hidden representation of states at each layer. We label our dataset according to whether at the given state the agent is facing a wall, an object, or it is free to move forwards. For each layer, we fit two linear probes on top of the hidden representations: one to predict if the agent is facing a wall and the other if it is facing an object.

### v2

@@ 3.1.1 State abstraction (optional)
where $\hat{P}^{g}$ is the empirical state transition function based on the abstract datasets ${\mathcal{D}}^{g}$, obtained in step 2. We then fine-tune a pre-trained large language model on ${\mathcal{D}}^{SFT}$ using the standard next-word prediction objective. We integrate description of the goal $g$ and the state $s^{g}_{0}$ into a prompt and set the sequence $[a_{0}^{*,g},\ldots,a_{n_{g}}^{*,g}]$ as the expected completion. We expect that the fine-tuned language model combined with the state abstraction function $\mathrm{LLM}^{abstrct}$ can effectively act as a proxy for the general, goal-conditioned policy $\pi^{*}$, generalizing to any new goal $g\notin{\mathcal{G}}^{tr}$ and previously unobserved low-level state $x\in{\mathcal{X}}$.
@@ 4 Related Work
LLMs for decision making. There is growing interest in using general-purpose LLMs directly as decision-making agents (Yao et al., 2023b). Various prompting techniques, such as chain-of-thought (CoT) (Wei et al., 2023) and self-reflection (Ji et al., 2023), have been developed to enhance LLMs’ abilities in long-term planning tasks. Yet, prompting alone is insufficient for solving complex tasks in dynamic environments (Szot et al., 2024; Finn, 2024). To effectively utilize the prior knowledge of LLMs, they must be grounded in the environment dynamics. This can be achieved either through in-context learning (Wang et al., 2023; Wu et al., 2023) or fine-tuning (Carta et al., 2023; Tan et al., 2024; Brohan et al., 2023a). A key limitation of in-context learning is its restricted window size. In this work, we focus on fine-tuning; however, unlike prior studies, we significantly reduce the requirements on the input data.
LLMs as data enhancers. In conventional goal-conditioned RL, datasets of state-action transitions require augmentation with goal-dependent rewards, often through human annotation or learning from demonstrations (Ziebart et al., 2008; Fu et al., 2018; Bahdanau et al., 2018). Recent studies show pre-trained LLMs can generate task-specific rewards (Yu et al., 2023b; Ma et al., 2023; Xie et al., 2023), though most rely on costly iterative prompting. We reduce these costs by approximating LLM-induced reward functions with proxy neural networks and assigning rewards in abstracted state spaces, lowering labeling needs. Similar to Peng et al. (2023), our approach uses natural language to guide the elimination of irrelevant state features.
Language-conditioned RL. Prior work on language-conditioned policies often assumes access to ground-truth rewards (Jiang et al., 2019; Co-Reyes et al., 2018), real-time experimentation (Fu et al., 2018; Bahdanau et al., 2018; Mirchandani et al., 2021), or expert demonstrations paired with language annotations (Stepputtis et al., 2020; Xiao et al., 2023; Brohan et al., 2023b, a). Our approach learns from offline, unlabeled datasets that are potentially suboptimal and lack reward signals. While most language-conditioned RL studies evaluate on synonymous commands seen during training (Lynch & Sermanet, 2021; Nair et al., 2022), we focus on novel instructions corresponding to previously unsolved goals which is a significantly more challenging setup that only a few related works attempt to address (Xiao et al., 2023; Brohan et al., 2023a; Jang et al., 2022).
See Appendix A for an extended discussion of related works.
@@ 5 Experiments
Table 1: Online evaluation of generalization performance. Results averaged over 400 $(g,s_{0}^{g})$ pairs
@@
DeepSeek-R1 (distilled to Llama-8B)a train/test train/test 32 (±3.6) 379 (±14.4) 40 (±0.5)
@@
  • a
    DeepSeek-R1 generates a reasoning trace per action, resulting in  100× longer responses by token count and inference time. Consequently, results are averaged over 50 $(g,s_{0}^{g})$ pairs.
Experimental Setup. We require a controlled with environment where a wide variety of distinct goals can be specified, whose state-spaces can be represented in natural language. We choose the BabyAI (Chevalier-Boisvert et al., 2018) environment as our main benchmark and the Webshop environment (Yao et al., 2023a) as an additional environment with an increased complexity of goals and actions, demonstrating the adaptability of TEDUO beyond grid-world environments. A detailed discussion on these choices can be found in Appendix A. BabyAI is a grid world platform for instruction following where an agent receives natural language goal instructions such as: “Go to the tile (3,2)” or “Look behind the green locked door”. The grids can consist of multiple rooms connected by open or locked doors and different distractor objects that the agent can interact with (boxes, keys, balls, etc.). The action space $\mathcal{A}$ consists of several navigation primitives (forward, pickup, etc). Webshop is a simulated e-commerce environment where an agent given product requirements must locate corresponding items by navigating a website. This environment is particularly relevant due to its dynamically evolving action space. The available actions vary by state due to changes in the website’s UI elements. Furthermore, the agent must generate linguistic inputs (e.g., search queries) to use the search bar. This highlights the necessity for the fine-tuned LLM to avoid catastrophic forgetting (Luo et al., 2025) and produce coherent keywords for narrowing searches. We work with the simplified HTML representation provided in the environment as our abstracted state space. Due to space constraints, results for the Webshop environment are presented in the Appendix B.1.
5.1 Q1: Online Evaluation: Generalization Benchmark
Setup. We choose the collection of Synth environments–the most complex environments not requiring state memory from BabyAI–as the main test bed for TEDUO. All environments are constructed as a 22x22 grid and containing 9 rooms. They differ in the type, position, and color of the distractors. The tasks include goals such as “go to the {color} {object}”, “pick up the {color} {object}”, or “put the {color} {object} next to the {color} {object}”. A list of 500 goals is used as ${\mathcal{G}}^{tr}$. The set of testing goals contains 100 goals that are semantically distinct from those in ${\mathcal{G}}^{tr}$. The set of testing goals is augmented by asking GPT-4 to paraphrase the original commands provided by BabyAI. We train a Llama-3-8B model with TEDUO based on a dataset ${\mathcal{D}}$ containing 800k non-unique state-action-state triplets generated according to a policy that is a random mixture of default policies from BabyAI [^3] Codebase to reproduce the main results of this paper can be found at this link. (see Appendix C.3 for details) .
Baselines. We compare our fine-tuned Llama-3-8B agent with non-fine-tuned LLMs: DeepSeek-R1 (DeepSeek-AI, 2025), Llama-3-8B and Llama-3-70B (Llama Team, 2024) using a) vanilla and b) CoT prompting (Wei et al., 2023) with additional demonstrations provided in-context (in-context+CoT). The latter integrates expert demonstrations generated during step 2 of TEDUO to test the in-context learning ability of the LLM. Following recent works (Mezghani et al., 2023; Li et al., 2022; Cao et al., 2023), we also compare against BabyAI-IL-bot, the baseline proposed by the authors of BabyAI (Chevalier-Boisvert et al., 2018), which is the combination of a GRU to encode the instruction, CNN+FILM layers to encode the grid and an LSTM memory. We train this method via imitation learning on the policy generated by TEDUO, steps 1&2. Implementation details of can be found in Appendix C.8.1.
@@ 5.1 Q1: Online Evaluation: Generalization Benchmark
Prior knowledge of LLMs is insufficient. We find that non-fine-tuned LLMs, irrespective of their parameter count or prompting method struggle in solving the seemingly simple tasks from the BabyAI environments. Low success rate and high invalid action ratios indicate the inability of LLMs to understand the dynamics of the environment. Common failures include only using the action “move forward” without considering the agent’s direction or attempting final actions (e.g. door opening) without first navigating to the correct location. This underscores the need for developing data-efficient methods for distilling knowledge of the environment-dynamics into LLMs. Our fine-tunning strategy brings the success rate from 17% to 65% in its training setting and 45% for testing on novel goals.
Unlocking generalization. We evaluate the generalization abilities of the fine-tuned TEDUO-Llama-3-8B model to new environments and goals. When tested on new environments unseen during training, a performance drop of 12% is observed, significantly smaller than BabyAI-IL-bot with a drop of 24%. This can be explained by the overfitting of the baseline due to the limited offline training data. TEDUO-Llama-3-8B benefits from the zero-shot capabilities of the pre-trained LLM. This effect is even more pronounced with new goals–TEDUO experiences only an 8% decrease in success rate, compared to the 40% drop for the BabyAI-IL-bot. Overall, TEDUO achieves nearly three times better performance than the RL baseline when generalizing to both new natural language commands and environments. Appendix B.6 shows success rates by goal type. Additional experiments demonstrating TEDUO’s generalization to larger, unseen BabyAI grids after training on smaller grids are presented in Appendix B.7. Performance remains robust for grids three times larger than those seen during training, confirming TEDUO’s generalization ability.
@@ 5.2 Q2: Online Evaluation: Ablation Study
Setup. With the same experimental setup, we compare our full fine-tuning pipeline with its ablations. After obtaining the abstract datasets ${\mathcal{D}}^{g}$ with the first step of TEDUO, we generate goal-conditioned policies with naive behavioral cloning (step 1 + GCBC). We also compare our fine-tuned Llama-3-8B against the performance of the GCRL policies obtained with offline Q-learning in step 2. Note, neither of GCBC nor GCRL can generalize to new, previously unseen language commands. Therefore, in this study, we are only looking at performance on goals from ${\mathcal{G}}^{tr}$. Ablation of the abstraction function is delayed to the next section.
Table 2: Ablation study. Results averaged over 400 $(g,s_{0}^{g})$ pairs.
@@
Results. The results of GCBC and GCRL can be seen as ablations of our pipeline. We first note that the success rate of naive behavioral cloning is low, indicating low fidelity of the data collection policy and highlighting the need for incorporating offline policy-learning methods. Moreover, the significantly improved performance of the Q-learning policies (GCRL) validates the effectiveness of the first two steps within our pipeline. The synthetically constructed MDPs are meaningful offline constructs that yield policies effective during online testing. Finally, the improved performance of our fine-tuned Llama-3-8B over the Q-learning/GCRL policies on training goals and environments confirms the importance of the third step of our method and suggests that the ungrounded, prior knowledge of large language models improves generalization to new previously unseen states.
@@ 5.3 Q3: Learning and exploiting core skills
We wish to investigate if by learning the optimal policies for diverse goals and environments, the LLM can integrate core skills required to achieve these goals and whether such skills can be transferred across tasks. We also investigate the aspect of skill compositionality. Does the prior knowledge of the LLM, now grounded in the environment dynamics, suffice to compose together learned skills?
(a) Type A
(b) Type B
(c) Type C
Figure 3: Three simple environment types.
@@
Setup. We are working with three types of simple environments illustrated in Figure 3. The position and color of the door and box vary across different instantiations of the environments. We use type A and B environments for training and type C environments for testing. We note that tasks from type C environments require the internalization of three core skills: moving to a given location, opening a door, picking up a box. The skill of moving to a location can be obtained from both environments A and B, but the skill of picking up a box or opening the door can only be obtained from one of the environments, A or B, respectively. This setup allows us to investigate the transferability of learned skills across environments and their compositionality.
Results. Table 3 shows that agents trained on one type of environments only fail to generalize. TEDUO A reaches a 99% success rate in tasks without closed doors (type A grids), but consistently fails when the goal is behind a door. TEDUO B achieves an 81% success rate in new grids from Type B but cannot generalize to Type C. The fact that TEDUO A&B can generalize to the environment C that requires a combination of both skills independently seen during training with a high success rate of 60% indicates that the fine-tuned LLM does not merely memorize optimal trajectories for individual tasks. Instead, it learns core, generalizable abilities that can be combined to solve novel tasks. This result emphasizes the significance of multi-skill learning for successful generalization that TEDUO enables.
Table 3: Performance on type C test tasks. TEDUO A and TEDUO B have been trained in only one environment whereas TEDUO A&B has been trained in both.
@@
One of the core skills required to successfully solve the tasks is to identify whether the agent at its location is facing an object or a wall, or it is free to move forwards. This section provides additional insights into the behavior and internal state representation of the LLM fine-tuned with TEDUO in comparison to a base LLM.
@@
Setup. As in the main evaluation benchmark, we operate within the Synth environments and generate a dataset with 10 random goals and 512 states per each goal. We embed each goal-state pair into the prompt template for eliciting actions and LLM fine-tuning and pass them through both the base and fine-tuned Llama-3-8B from experiments 5.1 and 5.2. We record the logprobabilities of the tokens [0, 1, …, 6] as well as the hidden representation of states at each layer. We label our dataset according to whether at the given state the agent is facing a wall or an object. For each layer, we fit two linear probes on top of the hidden representations: one for wall and one for object detection.